OpenStudy (anonymous):

expand log8(5v^2)(r^3 )

Nnesha (nnesha):

\[\huge\rm log_8(5v^2)(r^3)\] familiar with log property ?

OpenStudy (anonymous):

i have no idea what I'm supposed to do

Nnesha (nnesha):

quotient rule\[\huge\rm log_b y - \log_b x = \log_b \frac{ x }{ y}\] to condense you can change subtraction to division product rule \[\huge\rm log_b x + \log_b y = \log_b( x \times y )\] addition ----> multiplication power rule \[\huge\rm log_b x^y = y \log_b x\] you just need to know these properties of log that's it!

OpenStudy (anonymous):

so would the answer be 5log8+2logv+3logr ?

Nnesha (nnesha):

base would be the same for all log

Nnesha (nnesha):

and 5 isn't an exponent.

OpenStudy (anonymous):

log8+2log8+3log8, but where would the 5,r, and v go?

Nnesha (nnesha):

8 is base it would be like this \[\log_8\]

OpenStudy (anonymous):

would it be 2log8(5v)+3log8(r) assuming the 8s were bases

Nnesha (nnesha):

sorry the internet connection was disconnected